i need to evaluate/compute cos^2(pi/8)-sin^2(pi\8) thanks
Answer by lwsshak3(11628)

i need to evaluate/compute cos^2(pi/8)-sin^2(pi\8)
identity: cos(2x)=cos^2(x)-sin^2(x)
..
cos^2(pi/8)-sin^2(pi\8)=cos[2*π/8]=cos(π/4)=√2/2